OpenStudy (anonymous):

Factor and find all solutions of x^4 - 4x^3 - x^2 +16x = 12

OpenStudy (anonymous):

\[x^4-4x^3-x^2+16x=12\]

OpenStudy (anonymous):

I just don't know how to start it...

OpenStudy (anonymous):

Start with: \[ x^4-4x^3-x^2+16x-12=0 \]

OpenStudy (anonymous):

Yea but then i can't take out the GCF

OpenStudy (anonymous):

Because it was not a GCF to begin with.

OpenStudy (anonymous):

But that won't help you find roots.

OpenStudy (anonymous):

That's why i'm confused lol

OpenStudy (anonymous):

One think we can try is splitting the \(-x^2\) term.

OpenStudy (anonymous):

I was thinking maybe you could group them...

OpenStudy (anonymous):

But then how would you bring the 12 into it

OpenStudy (dan815):

hmm

OpenStudy (anonymous):

(By the way, I'm going of intuition here, this might not work, but it's an attempt)

OpenStudy (dan815):

u can guess a bunch but a nice way to see where ur zeros lie is to take first and 2nd derivatives

OpenStudy (dan815):

and graph and see where ur zeros should be

OpenStudy (anonymous):

I was never really taught how to graph these...

OpenStudy (dan815):

they give u nice ones, try guessing for this one though, 1,-1, 2,-2, 4,-4...

OpenStudy (anonymous):

Notice: \[ \color{red}{1}x^4\color{blue}{-4}x^2 +ax^2+\color{red}{b}x^2\color{blue}{+16}x-12 \]We want: \[ \frac{1}{-4} = \frac{b}{16} \]Which would imply that \(b=-4\).

OpenStudy (anonymous):

Now since \(a+b=-1\) and \(b=-4\), that means \(a+-4 = -1\implies a=3\).

OpenStudy (anonymous):

Where did u get the extra -x^4 from?

OpenStudy (anonymous):

This gives us: \[ (x^4-4x^3+3x^2)+(-4x^2+16x-12) \]

OpenStudy (anonymous):

What you could do is a binomial times a trinomial, no?

OpenStudy (anonymous):

Now, we can pull out the GFCs: \[ x^2(x^2-4x+3)+-4(x^2-4x+3) \]

OpenStudy (anonymous):

Which gives us: \[ (x^2-4)(x^2-4x+3) \]

OpenStudy (anonymous):

This method is called "splitting the middle"

OpenStudy (anonymous):

\[(x^4-4x^3)(-x^2+16x-12)\] Could u just do that from the very first equation?

OpenStudy (anonymous):

No, you can't. They are being added, not multiplied.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Besides, we aren't completely done yet.

OpenStudy (anonymous):

\[ x^2-4 = x^2-2^2 = (x-2)(x+2) \]This is called difference of squares.

OpenStudy (anonymous):

\[ x^2-4x+3 = (x^2-x) + (-3x+3) = x(x-1)+-3(x-1) = (x-3)(x-1) \]Splitting the middle another time.

OpenStudy (anonymous):

Sry i gtg i'll just try and ask my friend tommorrow morning u were a great help i just ran out of time x.x

OpenStudy (anonymous):

Final solution:\[ (x-2)(x+2)(x-3)(x-1)=0 \]